Field at End of Line of Charge A charged rod of length L = 5.60 m lies centered on the x axis as shown. The rod has a linear charge density which varies according to λ = ax where a =-583 C/m2 -L/2 +L/2 What is the total charge on the rod? 0.0000 C 4pts You are correct. Your receipt no. is 154-6850.) PreviousTries What is the x component of the electric field at a point on the x axis a distance of D = 7.80 m from the end of the rod? 4pts Submit Answer Tries 0/10
Answer #1

To calculate total charge on the rod and the electric field at P, we have to consider an elemental length dx on the rod at a distance 'x' from the left end of the rod (i.e. from origin) .

Given, the linear charge density of rod is

  lambda = ax (a =-583 ,,C/m 2 =-58.3 x 10-6 C/m-)

So, charge on the elemental length dx will be

  dq = lambda , dx = ax , dx

So, total charge on the rod will be

dq = I ax dx = aL2

Putting the values of a and L , we get

Q = rac{-58.3 imes 10^{-6} imes (5.60)^{2}}{2} = -9.14 imes 10^{-8} , , C

Hence, the total charge on rod is -9.14 x 10-8 C .

Electric Field at P :

Electric field due to a point charge 'q' at a distance 'r' can be calculated as

E = rac{1}{4 pi epsilon _{0}} , rac{q}{r^{2}}

Where,

  4TTEQ

But for this case we can not use this equation directly because the charge is not fixed at a point (or it is varying with x).

So, electric field at P due to the elemental charge da (= lambda , dx) will be

A dr 1 dq dE =

Where, r = L + D - x (as you can see in the figure)

So,   

  E = int_{0}^{L}dE = rac{1}{4 pi epsilon _{0}} , int_{0}^{L}rac{ax , dx}{(L + D -x)^{2}}

  = rac{a}{4 pi epsilon _{0}} , int_{0}^{L}rac{x , dx}{(L + D -x)^{2}}

  L+ D

L+D In (D) - In(L + D)-1 rl 4πέρ

  al In ( rl _- 4πέρ

Putting all the values , we get

5.60 5.607.807.80 7.80 -_58.3 × 10-6 × 9 × 109 × |ln (

-524.7 × 103 ×「一0.54 11+0.7179

ー-92.766 × 103 N/C (only in x - direction )

Te negative sign is indicating that the electric field is in negative x (i.e. -x ) direction.

NOTE :

The integration :

If

  (a - r)2

Let, a - x = t

or, x = a - t

or, dx = - dt

then,

  (a -t) dt t2 [(t - a) dt t2 dta dt t2

al

or,   I = ln , (a - x) + rac{a}{a - x} + C

For any doubt please comment and please give an up vote. Thank you.